Operation Summary for January 17, 2011 - January 19, 2011
Monday Jan. 17
The day shift began with Operations monitoring the Tevatron (TEV) store 8427
(181.88E30) and the Recycler stash (99.2E10), with the Antiproton Source (Pbar) stacking (15.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson
During the day shift, Operations monitored the accelerator complex.
During the evening shift, Operations terminated store 8427 at 9:19 PM. Operations established store 8428 at 10:42 PM with an average initial luminosity (AIL) of 312.22E30.
Tuesday Jan. 18
The midnight shift began with Operations monitoring store 8428 (228.62E30) and the stash (65.6E10), with Pbar stacking (15.7E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations reported that a network storm hit at 4 AM and halted the MiniBooNE beam until 4:30 AM. The rest of the complex ran through the storm, which finally cleared up about forty minutes after it started. Shortly after the network storm, intensity monitoring systems began reporting anomalous values and operators halted beam for 90 minutes while the error was diagnosed and understood to be a consequence of the network storm.
During the day shift, Operations rebooted the Booster Gradient Magnet Power Supply, which interrupted the Booster beam for about 20 minutes and then a Linac RF station (LRF1) tripped off and interrupted beam for another 20 minutes.During the evening shift, Operations terminated store 8428 at 5:30 PM. Operations established store 8429 at 6:42 PM with an AIL of 344.48E30.
Wednesday Jan. 19
The midnight shift began with Operations monitoring store 8429 (154.87E30) and the stash (148.4E10), with Pbar stacking (28.0E10), and with NuMI and MiniBooNE taking beam.During the midnight shift, Operations monitored the accelerator complex.
The Plans for Wednesday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 01/19/2011 email Fermilab|
|Security, Privacy, Legal|